OpenStudy (anonymous):

solve 4√x(only)+7=23

OpenStudy (anonymous):

\[4\sqrt{x}+7=23\] \[4\sqrt{x}=16\] \[\sqrt{x}=4\] \[x=4^2=16\]

OpenStudy (anonymous):

you isolate the radical when you are solving an equation. just means put it on one side of the equal sign.